Find a closed form of the recurrence given by an = -6+12-an-1; do = 2. Show work.
Answer #1

An = -6 + 12 an- no = 2 an =-6 + 12 an- -6+ 12 (-6 +12 an-2) 6 + 19(-6) +1440n-2 = -6+126-6) + 144 (-6+ 1201-3) = 6+ 12(-6) ++ 12 qn = (-6) +12(-6) + ... + (12 an: Set i = n, an = (-6) +126-6) +12+(6) + ... + (12) = (-6) + 12(-6) + 13C634 ...+(127.2